OpenStudy (anonymous):

question written better in the convo. is {2} an element of {3,2,1} 1022 {s| s=2^n-2,n element N} 3004 {x| x= 3n+1, n element N}

OpenStudy (anonymous):

is \[{2} \in {3,2,1} ? 1022 {s| s=2^n-2,n \in N} ? 3004 {x| x=3n+1, n \in N} \] ?

OpenStudy (aum):

Is 2 present in {3, 2, 1} ?

OpenStudy (anonymous):

yes.. so it is an element?

OpenStudy (aum):

Yes.

OpenStudy (aum):

For the next one, try various integer values in \(2^n-2\) and see if any of them give 1022.

OpenStudy (anonymous):

sorry but im lost in tht

OpenStudy (aum):

{s| s=2^n-2,n element N} n is a natural number (belongs to set N). That means n = 1, 2, 3, 4, ...... n = 1, s = 2^1 - 2 = 0 n = 2, s = 2^2 - 2 = 2 n = 3, s = 2^3 - 2 = 6 n = 4, s = 2^4 - 2 = 14 .... So instead of saying the set is: {0, 2, 6, 14, .......} the set is expressed by a formula: {s| s=2^n-2,n element N}. So keep trying higher n values to see if 1022 is an element of the set.

OpenStudy (anonymous):

ooh ok let me try

OpenStudy (aum):

Don't have to try in sequence. You can skip a few and if it exceeds 1022, step down a little bit.

OpenStudy (anonymous):

2^10 -2

OpenStudy (aum):

Yes, when n = 10, it gives 1022 and therefore 1022 is an element of the set.

OpenStudy (anonymous):

so it is an element?

OpenStudy (aum):

Yes. How about the last one?

OpenStudy (anonymous):

it is an element..... 3x1001+1

OpenStudy (anonymous):

1 more quest please

OpenStudy (anonymous):

are they all \[\subset or are there \subseteq\]

OpenStudy (aum):

not sure what the question is.

OpenStudy (anonymous):

ok are they all subsets or are they proper subsets

OpenStudy (aum):

If it is about set notations then: http://www.purplemath.com/modules/setnotn.htm

OpenStudy (anonymous):

ok so they are all subsets because in a proper subset .. there is at least one elemnt of A that is not in B

OpenStudy (anonymous):

@aum

OpenStudy (aum):

Sorry, the notification system on this site has not been working properly lately and I did not know until now you had posted a reply. {2} and {3, 2, 1} The first set is completely contained in the second set AND the first set is NOT equal to the second set. Therefore, it is a proper subset. Same answer for the other two as well if you consider {1022} as one set and {s| s=2^n-2,n element N} as the other set. The first set is completely contained in the second set and the two sets are not equal. Therefore, it is a proper subset. {3004} is a proper subset of {x| x= 3n+1, n element N}